# Can Test-time Computation Mitigate Memorization Bias in Neural Symbolic Regression?

### v1

@@ 4.1 Method
For evaluation, we constructed two test datasets: not_included and baseline, each containing $150$ expressions. For the not_included set, we removed every $e_{\mathrm{templ}}$ appearing in the training data. In contrast, the baseline set was sampled directly from the generator $p_{\mathcal{E}}$ without any filtering. We associated each expression with $100$ data points, generated in the same way as during training. We set the beam size to $5$ for this experiment.
@@
Note that these $m$ evaluation points are distinct from the inputs provided to the model at test time. By definition, $R^{2}\leq 1$, and values closer to $1$ indicate that the predicted outputs closely match the true equation. In our experiments, we counted the number of predictions whose $R^{2}$ exceeds thresholds of $0.5,0.9,0.95,0.99,0.999,0.9999$, and $0.9999$, respectively. This allows us to assess the model’s ability to fit the data under both moderate and stringent accuracy requirements.
4.2 Results
Figure 1: NeSymReS exhibits strong fitting performance on the baseline dataset but performs poorly on expressions from the not_included dataset, whose tree structures are absent from the training data.
We first evaluated NeSymReS’s ability to generate novel expressions using the not_included dataset. As this dataset comprises instances of $e_{\mathrm{templ}}$ unseen in the training data, the model is expected to produce previously unseen tree structures. However, only 2.90% of the generated expressions has tree structures (excluding constants) that are not present in the training set. This result indicates strong memorization bias in NeSymReS and reveals the constrained size of its search space.
@@ 4.2 Results
Next, Figure 1 illustrates NeSymReS’s fitting performance on the not_included and baseline datasets. The results indicate a substantial drop in performance for expressions whose tree structures are not present in the training data, compared with those sampled randomly. This suggests that for expressions not seen during training, the model’s memorization bias directly leads to poor numerical accuracy. This result also helps explain why NSR methods often fail to achieve high performance on expressions with many input variables; an increase in the number of input variables leads to an expanded search space, thereby increasing the likelihood that a given expression is absent from the training set.
4.3 Theoretical Analysis
The preceding results show that NeSymReS struggles to generate expressions that were not encountered during training. In this section, we examine the theoretical basis of this limitation. We show that Transformers inherently lack the capacity to generate expressions compositionally while accounting for their numerical characteristics. We introduce a simplified version of the symbolic regression task and show that Transformers are not expressive enough to solve the task.
We define the last-token prediction problem as the task of predicting the most suitable final token in an otherwise complete mathematical expression. Although predicting the entire optimal expression is NP-hard [33], this task is much easier since the search space is limited to several leaf tokens. We present a formal definition of this task in the following.
We first introduce $\mathrm{expr}:\Gamma^{\ast}\times\mathbb{R}^{(d+1)*n}\!\longrightarrow\!\mathcal{E}$, a function that maps a token sequence $\mathbf{s}$ to the most appropriate expression $e_{\mathbf{s}}$ that can be represented by $\mathbf{s}$, taking numerical data $\mathcal{D}$ into account. Since the token sequence may contain the placeholder token $C$ representing constants, the mapping is tasked with identifying the optimal values for these constants and transforming the sequence into a corresponding expression tree.
Definition 1 (Last-token prediction problem).
Given numerical data $\mathcal{D}$ of $n$ features-value pairs $(\mathbf{x}_{i},y_{i})\in\mathbb{R}^{d}\times\mathbb{R}$, a metric $\mathcal{L}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$, and an incomplete token sequence $\widetilde{\mathbf{s}}$ that forms a prefix representation of an expression when terminated with a leaf token, the last-token prediction problem asks for finding a leaf token $u^{\ast}$ such that:
For the analysis, we assume a bounded-depth log-precision Transformer as in [8, 19, 18, 24], a realistic setting where the intermediate computation values of the Transformer are limited to $O(\log k)$ bit precision, with $k$ denoting the maximal length of the input sequence. We now present the theoretical result stating that Transformers with bounded size cannot solve the last-token prediction problem.
Theorem 1.
Assume $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$. For any integer $D$ and any polynomial $Q$, there exists a problem size $m$ such that no log-precision Transformer defined in Section E.1 with depth $D$ and hidden dimension $d\leq Q(m)$ can solve $\operatorname{LastTokenPrediction}(m)$.
We show the above theorem by leveraging circuit complexity theory. Specifically, $\mathsf{TC}^{0}$ and $\mathsf{NC}^{1}$ are types of circuit complexity classes, and it is generally conjectured that $\mathsf{TC}^{0}\subsetneq\mathsf{NC}^{1}$. Prior work [19] shows that log-precision Transformers can be simulated with $\mathsf{TC}^{0}$ circuits. We provide a proof for the above theorem by showing that the complexity of the last-token prediction problem is lower bounded by $\mathsf{NC}^{1}$. Detailed specifications of the problem setting and proof of the theorem are provided in Appendix E.
Although the final token of a mathematical expression is arguably the easiest to predict among its components, the above theorem shows that even this seemingly simple task presents substantial difficulties for Transformer models. The reason why a Transformer fails to solve this task is that it cannot numerically verify the expressions it has generated so far. This limitation helps explain why NeSymReS tends to memorize and reproduce familiar expressions rather than generate novel ones compositionally.
5 Can Test-time Strategies Mitigate Memorization Bias?
@@
The results from the previous section indicate that the search space of NeSymReS is confined to expressions seen during training. We also analyzed that the reason lies in the Transformer’s lack of ability to compositionally generate predictions. Our hypothesis is that providing the model with hints about which tokens are appropriate could help overcome memorization bias. In this section, we explore how different test-time strategies affect memorization bias and numerical accuracy in neural symbolic regression. We begin by introducing the three test-time strategies employed in our experiments. The detailed explanation for the strategies are presented in Section B.
@@ 5.1 Test-time Strategies
Beam search serves as the default decoding strategy employed by NeSymReS. During decoding, Given a beam size of $b$, the decoding process generates $b$ candidate sequences via beam search. Each candidate’s constant placeholders are subsequently optimized using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [9]. The expression exhibiting the highest numerical accuracy on the test data is then selected as the model’s output. While the experiments in the previous section used a beam size of $b=5$, in this section we conducted experiments with a larger beam size of $b=150$. Since increasing the beam size does not provide the model with any additional information, our hypothesis is that simply adopting a decoding strategy with a larger beam size will not alleviate memorization bias.
Transformer-based planning for symbolic regression (TPSR).
TPSR [23] is a method that leverages MCTS during decoding time. In TPSR, the process starts by preparing a pre-trained NSR model (e.g., the NeSymReS model). Instead of relying on standard decoding methods like beam search, the method generates tokens using MCTS, where both the expansion and evaluation stages of MCTS leverage the pre-trained NSR model. In the expansion phase, to avoid unnecessary exploration, the set of expandable tokens is restricted to the top-$k_{\mathrm{max}}$ candidates based on the logits from the NSR model. During the evaluation phase, the NSR model first completes the remainder of the expression following the expanded token. The completed expression is then evaluated primarily based on its fitting accuracy, with additional consideration given to its complexity. In the experiments presented in this section, we used the default hyperparameter settings of TPSR as specified in the original paper; we set the number of rollouts to $r=3$, the number of expandable tokens to $k_{\mathrm{max}}=3$, and beam size for expression completion to $b=1$.
@@
Figure 2: Overview of NSR-gvs’s inference procedure. We first sample subtrees from the candidate pool, then supply them to the model together with numerical data. Then, the generated prediction is numerically verified and the self-verification feedback is used to update the candidate pool. This procedure is performed repeatedly to generate better predictions over time.
TPSR provides feedback to the model by assigning a reward to each token, reflecting the quality or appropriateness of that token. In contrast, we hypothesize that incorporating feedback at the subtree level as well may have a positive effect on the model. To this end, we propose NSR-gvs, a method grounded in the following intuition: expressions that fit the same numerical data well are likely to share common substructures.
We first train a slightly modified version of the NeSymReS model, where the model takes subtrees as prompts and generates expressions that incorporate them. We achieve this by extracting subtrees from the ground-truth expressions and feeding them to the model together with numerical data during training. Figure 2 illustrates the inference procedure of NSR-gvs. We generate multiple predictions iteratively by augmenting the model with varying prompts. For each iteration, we first sample subtrees from a pool of candidate subtrees, which were extracted from high-performing expressions in previous predictions. To maintain output diversity, we also occasionally sample subtrees from a random distribution. We then provide the sampled subtrees to the pre-trained model as prompts, along with the numerical data. After the model generates a prediction, it is automatically verified according to the fitting accuracy on the test data. Finally, the pool of candidate subtrees are updated based on the results of self-verification. This method can be formulated within the framework of reinforcement learning, and we provide a more detailed explanation in the appendix.
@@ 5.2 Results in Controlled Settings
First, we evaluate the impact of each test-time strategy on memorization bias and numerical accuracy in a controlled experimental setting. The experimental setup closely follows that described in Section 4. We train a NeSymReS model and a prompt-augmented model for NSR-gvs with the same training dataset for the same number of epochs. We evaluated the strategies using the not_included dataset, where we used the $R^{2}$ metric to evaluate numerical accuracy, and the number of novel expressions to evaluate memorization bias.
@@
Figure 3 shows how the test-time strategies perform under the controlled setting. In terms of the ability to generate novel expressions, TPSR, NSR-gvs, and their combination demonstrate strong performance. These results imply that strategies involving the provision of additional information during inference are more effective in reducing memorization bias. However, the result shows that high novelty in generated expressions does not necessarily imply high numerical accuracy. In some cases, acquiring the ability to generate novel expressions leads to a decrease in numerical accuracy (TPSR), whereas some strategies can improve numerical accuracy despite high memorization bias (large beam size).
5.3 Numerical Accuracy in Practical Settings
We additionally evaluate and compare the numerical performance of the test-time strategies under conditions that better reflect practical applications. A total of $10$ million expressions were used to construct the training dataset, employing all operators described in Section A without any restriction on operator types. We trained both a NeSymReS model and a prompt-augmented model on this dataset for $10$ epochs. For the test datasets, we prepared the following two sets:
  • AI-Feynman. This dataset consists of $91$ equations with up to five independent variables, extracted from the AIFeynman database [28]. It is commonly used in various studies to assess the performance of symbolic regression methods.
  • only_five_variables_nc. This dataset consists of expressions containing exactly five independent variables, making it a challenging dataset. The “nc” designation indicates that the expressions do not include constants, which simplifies the problems slightly; however, it remains more difficult than the first dataset. The dataset was constructed by sampling expressions from $p_{\mathcal{E}}$, filtering for expressions that include exactly five variables, and finally deleting its constants. This dataset is derived from the study of NSRwH [3], and we use the first $100$ expressions for evaluation.
Figure 4 demonstrates how the different test-time strategies perform under more practical settings. TPSR relatively performs slightly better than in the controlled setting; however, the general pattern of numerical accuracy remains consistent. These results demonstrate that, even in practical settings, test-time strategies that mitigate memorization bias do not always result in better performance.
Figure 4: Comparison of test-time strategies under practical settings. The figure on the left shows the performance on the AI-Feynman dataset, and the figure on the right presents results on the only_five_variables_nc dataset.
5.4 Discussion
The experimental results show that while methods such as TPSR, which mitigate memorization bias, can lead to a drop in numerical accuracy, decoding with large beam size improves numerical accuracy despite retaining a high level of memorization bias. In this subsection, we discuss possible reasons why such phenomena occur.
Given that the additional information in TPSR and NSR-gvs is derived from self-verification, it should offer better alternatives beyond the model’s own logits, and is expected to assist in generating better expressions. Despite providing such useful information, the methods often underperform in terms of numerical accuracy compared to the simple strategy of increasing beam size. This suggests that the Transformer struggles to leverage the additional information effectively, and in some cases, it might even be negatively impacted by it. For example, when the Transformer encounters an unfamiliar prefix within an partially constructed expression, it may become confused and could complete the expression with suboptimal tokens.
We therefore argue that providing additional information at test time in a way that is easy for the Transformer to leverage is important for developing a truly generalizable NSR approach.
@@ 6 Conclusion
In this work, we first identified a major drawback of standard NSR models; they can hardly generate novel expressions that were not seen during training, highly limiting the search space. We also provide theoretical analysis to explain this memorization bias and showed that Transformers are incapable of generating expressions in a compositional way, while taking numerical data into account. We then demonstrated that incorporating additional information to the model during test-time can reduce memorization bias. However, we also show that mitigating memorization bias does not necessarily lead to higher numerical accuracy. The main limitation of this work is the absence of a method that simultaneously mitigates memorization bias and improves numerical accuracy to a significant extent. In future work, we aim to build on the findings of this study to design symbolic regression methods with improved generalizability.
@@ Appendix A Details for NeSymReS
In the first step for generating the expression $e^{\ast}$, the unary-binary tree structure $T$ is generated randomly within the limits of a maximum depth of $6$. In the third step, the total number of constants added to the expression is also limited to a maximum of $6$. The binary and unary operators $\mathcal{O}_{\mathrm{bin}}\cup\mathcal{O}_{\mathrm{un}}$ are shown in Table 2. Other hyparparameters are specified in Table 3, where $\mathcal{LU}$ denotes the log-uniform distribution.
@@
$d$ Dimension for input variables $5$
@@
The NeSymReS model consists of two architectural components: the numerical encoder $enc_{\mathrm{num}}$ and a decoder $dec$. The numerical encoder processes the numerical data $\mathcal{D}$, represented as a tensor of shape $(b,n,d)$, where $b$ denotes the batch size, $n$ the number of input points, and $d$ the sum of dependent and independent variables. First, an embedding layer converts the numerical data into a higher dimensional tensor $\mathcal{D}^{\prime}$ of shape $(b,n,h)$. This tensor is then processed by a $5$-layer set-transformer [15] encoder that outputs a new tensor $Z_{num}$ of shape $(b,s,h)$, where $s$ denotes the number of embedding vectors produced by the encoder. The resulting tensor $Z_{num}$ is subsequently passed to the decoder $dec$, a five-layer standard Transformer decoder that auto-regressively generates the corresponding expression token by token. We set $b=200$, $h=512$, and $s=32$ for our experiments.
@@
During training, cross-entropy loss is used as the objective function, and teacher forcing [26] is applied during next-token prediction. The AdamW [17] optimizer is employed with an initial learning rate of $10^{-4}$. After $4000$ steps, the learning rate is adjusted proportionally to the inverse square root of the number of steps taken.

### v2

@@ 5.1 Reproduction Bias in Simplified Setting
For evaluation, we constructed two test datasets: not_included and baseline, each containing $150$ expressions. For the not_included set, we removed all $e_{\mathrm{templ}}$ appearing in the training data. In contrast, the baseline set was sampled directly from the generator $p_{\mathcal{E}}$ without any filtering. We associated each expression with $100$ data points, generated in the same way as during training. We set the beam size to $5$ for this experiment.
@@
Note that these $m$ evaluation points are distinct from the inputs provided to the model at the test time. By definition, $R^{2}\leq 1$, and values closer to $1$ indicate that the predicted outputs closely match the true equation. In our experiments, we counted the number of predictions whose $R^{2}$ exceeds the thresholds of $0.5,0.9,0.95,0.99,0.999,0.9999$, and $0.9999$, respectively. This allows us to assess the model’s ability to fit the data under both moderate and stringent accuracy requirements.
Result.
@@
Figure 1 shows the results for NeSymReS under the simplified setting. The left figure demonstrates NeSymReS’s ability to generate novel expressions using the not_included dataset. As this dataset comprises instances of $e_{\mathrm{templ}}$ unseen in the training data, the model is expected to produce previously unseen tree structures. However, the result indicates that NeSymReS struggles to generate expression trees beyond the training data across varying epochs. After $1000$ epochs of training, over 97% of the generated expression trees were direct copies from the training data, highlighting strong reproduction bias and revealing that the search space of NeSymReS is severely restricted. The right figure demonstrates how this reproduction bias negatively affects numerical accuracy, where NeSymReS’s fitting performance on the not_included and baseline datasets is compared. The results indicate a substantial drop in performance for expressions whose tree structures are not present in the training data, compared with those sampled randomly. This suggests that for expressions not seen during training, the model’s reproduction bias directly leads to poor numerical accuracy. This result also helps explain why NSR methods often fail to achieve high performance on expressions with many input variables; an increase in the number of input variables leads to an expanded search space, thereby increasing the likelihood that a given expression is absent from the training set.
Figure 2: Reproduction bias in transformer4sr with varying training dataset sizes. While small training dataset sizes (100K, 200K) exhibit stronger reproduction bias, scaling the training dataset size does not necessarily mitigate reproduction bias after a certain limit.
5.2 Reproduction Bias in Practical Setting
Due to the complicated training procedure of NeSymReS, the above analysis was carried out in a simplified setting. To examine whether reproduction bias is a general phenomenon, we conducted an additional analysis in a more practical setting using transformer4sr (Lalande et al., 2023), a method similar to NeSymReS but with a simpler training process. In transformer4sr, no dynamic transformations of expressions are applied during training, which makes it much easier than in NeSymReS to verify whether the generated expressions are included in the training data. We were also able to analyze the novelty of the expressions by explicitly taking into account the positions of the constant placeholder tokens. Consequently, the definition of reproduction bias slightly differs from that described in Section 5.1, and we therefore provide additional clarification in Appendix D.
We followed the model architecture, number of epochs, operator selection, and inference strategies described in Lalande et al. (2023). We constructed training datasets with varying size, with the largest one consisting of 1.51.5M expressions based on the practical setting used in the original paper. We used the full set of operator tokens, which are add, mul, cos, log, exp, neg, inv, sqrt, sq (squared), cb (cubed), and the number of independent variables was $6$. We constructed a test set similar to not_included in the previous analysis, which consists of $300$ expressions that were not included in the training data.
Figure 2 shows the result for transformer4sr’s ability to generate expressions beyond the training data. Even in a practical setting with 1.51.5M training expressions, less than $12\%$ of the expressions generated by transformer4sr were novel expressions (taking into account the position of the constant placeholder tokens) beyond the training data, and less than $6\%$ of the expressions had novel tree structures (excluding constants). The result also demonstrates that the effect of scaling training data on mitigating reproduction bias tends to saturate, suggesting that reproduction bias is likely to persist even when the training data size is further increased.
6 Can Test-time Strategies Mitigate Reproduction Bias?
@@
The results from the previous section indicate that the search space of NeSymReS is mostly confined to expressions seen during training due to reproduction bias. Since our theoretical analysis indicates that naively performing next-token prediction makes it difficult to generate novel expressions in a compositional manner, we investigated the possibility of devising inference-time computational techniques to reduce reproduction bias in this section. Our hypothesis is that providing the model with hints about which tokens are appropriate could help steer the model to generate expressions that were not included in the training data. We begin by briefly introducing the three test-time strategies employed in our experiments. The detailed explanation for the strategies is presented in Appendix B.
@@ 6.1 Test-time Strategies
The beam search serves as the default decoding strategy employed by NeSymReS. During decoding, given a beam size of $b$, the decoding process generates $b$ candidate sequences via beam search. The constant placeholders of each candidate are subsequently optimized using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm (Fletcher, 2000). The expression exhibiting the highest numerical accuracy on the test data is then selected as the model’s output. While the experiments in Section 5.1 used a beam size of $b=5$, in this section we conducted experiments with a larger beam size of $b=150$. Since increasing the beam size does not provide the model with any additional information, our hypothesis is that simply adopting a decoding strategy with a larger beam size will not alleviate reproduction bias.
Figure 3: Overview of NSR-gvs’s inference procedure. We first sampled subtrees from the candidate pool, then supplied them to the model together with numerical data. Then, the generated prediction is numerically verified and the self-verification feedback is used to update the candidate pool. This procedure is performed repeatedly to generate better predictions over time.
Incorporating MCTS.
TPSR (Shojaee et al., 2023) is a method that leverages MCTS during decoding time. In TPSR, the process starts by preparing a pre-trained NSR model (e.g., the NeSymReS model). Instead of relying on standard decoding methods like beam search, the method generates tokens using MCTS, where both the expansion and evaluation stages of MCTS leverage the pre-trained NSR model. In the expansion phase, to avoid unnecessary exploration, the set of expandable tokens is restricted to the top-$k_{\mathrm{max}}$ candidates based on the logits from the NSR model. During the evaluation phase, the NSR model first completes the remainder of the expression following the expanded token. The completed expression is then evaluated primarily based on its fitting accuracy, with additional consideration given to its complexity. In the experiments presented in this section, we used the default hyperparameter settings of TPSR as specified in the original paper; we set the number of rollouts to $r=3$, the number of expandable tokens to $k_{\mathrm{max}}=3$, and the beam size for expression completion to $b=1$.
@@
TPSR provides feedback to the model by assigning a reward to each token, reflecting the quality or appropriateness of that token. In contrast, we hypothesized that incorporating feedback at the subtree level as well may have a positive effect on the model. To this end, we propose NSR-gvs, a method grounded in the following intuition: expressions that fit the same numerical data well are likely to share common substructures.
We first trained a slightly modified version of the NeSymReS model, where the model takes subtrees as prompts and generates expressions that incorporate them. We achieved this by extracting subtrees from the ground-truth expressions and feeding them to the model together with numerical data during training. Figure 3 illustrates the inference procedure of NSR-gvs. We generated multiple predictions iteratively by augmenting the model with varying prompts. For each iteration, we first sampled subtrees from a pool of candidate subtrees, which were extracted from high-performing expressions in previous predictions. To maintain output diversity, we also occasionally sampled subtrees from a random distribution. We then provided the sampled subtrees to the pre-trained model as prompts, along with the numerical data. After the model generates a prediction, it is automatically verified according to the fitting accuracy on the test data. Finally, the pool of candidate subtrees are updated based on the results of self-verification. This method can be formulated within the framework of reinforcement learning, and we provide a more detailed explanation in Appendix B.
@@ 6.2 Results
We evaluated the impact of each test-time strategy on reproduction bias and numerical accuracy in an experimental setting. The experimental setup closely follows that described in Section 5.1. We trained a NeSymReS model and a prompt-augmented model for NSR-gvs with the same training dataset for the same number of epochs. We evaluated the strategies using the not_included dataset, where we used the $R^{2}$ metric to evaluate numerical accuracy, and the number of novel expressions to evaluate reproduction bias.
@@
Figure 4 shows how the test-time strategies perform under the simplified setting. In terms of the ability to generate novel expressions, TPSR, NSR-gvs, and their combination demonstrate strong performance. These results imply that strategies involving the provision of additional information during inference (TPSR and NSR-gvs) are more effective in reducing reproduction bias. However, the result shows that high novelty in generated expressions does not necessarily imply high numerical accuracy. In some cases, acquiring the ability to generate novel expressions leads to a decrease in numerical accuracy (TPSR), whereas some strategies can improve numerical accuracy despite high reproduction bias (large beam size). This suggests that, in some cases, Transformers struggle to leverage the additional information effectively. We therefore argue that providing additional information at test time in a way that is easy for the Transformer to leverage is important for developing a truly generalizable NSR approach. Viewed in this way, the use of subtrees at inference, as in the proposed method NSR-gvs, can be seen as a potentially valuable approach, since it contributes to mitigating reproduction bias and improving numerical accuracy.
@@ 7 Conclusion
In this work, we identified a major drawback of standard NSR models both theoretically and empirically. Our theoretical analysis shows that Transformers are incapable of generating expressions in a compositional way, while taking numerical data into account. We then examined the strategies that Transformers actually employ to generate expressions, and the results suggest that they mostly generate expressions copied from the training data, highly limiting the search space. Finally, we demonstrate that incorporating additional information to the model during test-time can reduce reproduction bias. In future work, we aim to build on the findings of this study to design symbolic regression methods that further improve generalizability.
@@ Appendix A Details for NeSymReS
In the first step for generating the expression $e^{\ast}$, the unary-binary tree structure $T$ is generated randomly within the limits of a maximum depth of $6$. In the third step, the total number of constants added to the expression is also limited to a maximum of $6$. The binary and unary operators $\mathcal{O}_{\mathrm{bin}}\cup\mathcal{O}_{\mathrm{un}}$ are shown in Table 2. Other hyparparameters are specified in Table 3, where $\mathcal{LU}$ denotes the log-uniform distribution. After the expression is sampled, the symbolic manipulation library Sympy (Meurer et al., 2017) is used to simplify any redundancy in the expression.
@@
$d$ Dimension for input variables 55
@@
The NeSymReS model consists of two architectural components: the numerical encoder $enc_{\mathrm{num}}$ and a decoder decdec. The numerical encoder processes the numerical data $\mathcal{D}$, represented as a tensor of shape $(b,n,d)$, where $b$ denotes the batch size, $n$ the number of input points, and $d$ the sum of dependent and independent variables. First, an embedding layer converts the numerical data into a higher dimensional tensor $\mathcal{D}^{\prime}$ of shape $(b,n,h)$. This tensor is then processed by a $5$-layer set-transformer (Lee et al., 2019) encoder that outputs a new tensor $Z_{num}$ of shape $(b,s,h)$, where $s$ denotes the number of embedding vectors produced by the encoder. The resulting tensor $Z_{num}$ is subsequently passed to the decoder decdec, a five-layer standard Transformer decoder that auto-regressively generates the corresponding expression token by token. We set $b=200$, $h=512$, and $s=32$ for our experiments.
@@
During training, cross-entropy loss is used as the objective function, and teacher forcing (Sutskever et al., 2014) is applied during next-token prediction. The AdamW (Loshchilov and Hutter, 2017) optimizer is employed with an initial learning rate of $10^{-4}$. After $4000$ steps, the learning rate is adjusted proportionally to the inverse square root of the number of steps taken.